OpenStudy (anonymous):

differentiate y= (e^x)/(1+x)

OpenStudy (anonymous):

please show step by step

OpenStudy (anonymous):

quotient rule \[\frac{d}{dx}\frac{f}{g}=\frac{gf'-f'g}{g^2}\]}

OpenStudy (anonymous):

Use the quotient rule. Which says: \[\frac{d}{dx}[\frac{f(x)}{g(x)}]=\frac{f'(x)g(x)-f(x)g'(x)}{(g(x))^2}\]. Let e^x=f(x) and (1+x)=g(x) So f'(x)=e^x g'(x)=1 Plug them in: \[\frac{(e^x)(1+x)-e^x(1)}{(1+x)^2}\]. Then you can simplify it. Tell me if you need more explanation :P

OpenStudy (anonymous):

so if you really want to do it sept by step

OpenStudy (anonymous):

what maleviolence19 said

OpenStudy (anonymous):

yes thanks guys

OpenStudy (anonymous):

malevolence sorry

OpenStudy (anonymous):

No problem rmalik :P

OpenStudy (anonymous):

so the answer i got would be \[ex^x/(1+x)2\]

OpenStudy (anonymous):

would that be right malevolence19 or satellite73